OpenStudy (anonymous):

Differentiate f(x)=ln(x+sqrt(x^2-1))

OpenStudy (anonymous):

\[\frac{d}{dx}\ln(f(x))=\frac{f'(x)}{f(x)}\] use \[f(x)=x+\sqrt{x^2-1}, f'(x)=1+\frac{x}{\sqrt{x^2-1}}\]